the problem is to find the slope of the line. (x/2)+(y/3)=1
i dont know if the slope is -1 or (-3x/6)
i think it is -1 but my sister said im wrong!!!! please help im so confused

Answer by scott8148(6628)

subtracting x/2 ___ y/3 = -x/2 + 1

multiplying by 3 ___ y = (-3/2)x + 3 ___ this is the slope-intercept form, y = mx + b

the slope is the coefficient of the x-term; in this case, -3/2